A company with name XYZ starts a new project with life 4 years. The company will put 20.000.000 USD in the project. The estimated sales per year is 1200 units a)Selling price for first year is 45.000USD/unit b)Variable cost for first yaer is 35.000USD/unit c)Fixed cost for first year is 2.000.000 USD d)Depreciation rates per year:20% , 32%,19.20%, 11.52% e)Net working capital requirements for the project are 20% of the projected annual sales at the beggining of each yea.r f) After first year, price of product, variable and fixed costs will increase with inflation glA the end of the 4 year-project, the equipment wil have 800.000 USD market value We also know the cost of capital which is 12% and the tax rate which is 35% What is the NPV, IRR and payback of this project?
Answer #1
ASSUME ANNUAL INFLATION =4%=0.04
a Selling Price per unit year1 $45,000
b=a*1.04 Selling Price per unit year2 $46,800
c=b*1.04 Selling Price per unit year3 $48,672
d=c*1.04 Selling Price per unit year4 $50,619
e Variable cost per unit year1 $35,000
f=e*1.0 Variable cost per unit year2 $36,400
g=f*1.04 Variable cost per unit year3 $37,856
h=g*1.04 Variable cost per unit year4 $39,370
i Fixed cost per unit year1 $2,000,000
j=i*1.04 Fixed cost per unit year2 $2,080,000
k=j*1.04 Fixed cost per unit year3 $2,163,200
l=k*1.04 Fixed cost per unit year4 $2,249,728
Present Value (PV) of Cash Flow:
(Cash Flow)/((1+i)^N)
i=Discount Rate=Cost of capital=12%=0.12
N=Year of Cash Flow
N Year 0 1 2 3 4
A Unit sales                1,200                 1,200               1,200                  1,200
B Selling Price per unit $45,000 $46,800 $48,672 $50,619
X=A*B Sales Revenue $54,000,000 $56,160,000 $58,406,400 $60,742,656
C Variable cost per unit $35,000 $36,400 $37,856 $39,370
D=B-C Unit Contribution Margin $10,000 $10,400 $10,816 $11,249
E=A*D TotalContribution $12,000,000 $12,480,000 $12,979,200 $13,498,368
F Fixed costs $2,000,000 $2,080,000 $2,163,200 $2,249,728
G=E-F Pretax profit (Before depreciation) $10,000,000 $10,400,000 $10,816,000 $11,248,640
H Depreciation Rate 20% 32% 19.20% 11.52%
I=H*$20million Annual depreciation $4,000,000 $6,400,000 $3,840,000 $2,304,000
J Accumulated depreciation $4,000,000 $10,400,000 $14,240,000 $16,544,000
K End of year book value $3,456,000 (20000000-1654400)
L=G-I Pretax profit (After depreciation) $6,000,000 $4,000,000 $6,976,000 $8,944,640
M=0.35*L Tax expenses $2,100,000 $1,400,000 $2,441,600 $3,130,624
P=L-M After tax profit $3,900,000 $2,600,000 $4,534,400 $5,814,016
Q=P+I After tax Operating Cash Flow $7,900,000 $9,000,000 $8,374,400 $8,118,016
R Investment cash flow ($20,000,000)
S Working Capital Need(20% of next year Sales Revenue) $10,800,000 $11,232,000 $11,681,280 $12,148,531
T Cash Flow due to change in working capital ($10,800,000) ($432,000) ($449,280) ($467,251) $12,148,531
U Pretax Salvage Value $800,000
V=U-K Gain on salvage ($2,656,000)
W=V*0.35 Tax on gain in salvage ($929,600)
Y=U-W After tax cash flow on salvage $1,729,600
Z=Q+R+T+Y NET CASH FLOW ($30,800,000) $7,468,000 $8,550,720 $7,907,149 $21,996,147
CUMULATIVE NET CASH FLOW ($30,800,000) ($23,332,000) ($14,781,280) ($6,874,131) $15,122,016
Payback Period=Period at which cumulative cash flow=NIL
PAYBACK PERIOD=4+(6874131/21996147)                           4.31 YEARS
SUM
PV=Z/(1.12^N) PRESENT VALUE OF NET CASH FLOW $       (30,800,000) $    6,667,857 $    6,816,582 $ 5,628,152 $   13,978,949 $ 2,291,540
NET PRESENT VALUE $            2,291,540
IRR(INTERNAL RATE OF RETURN) 14.91% (Using IRR function of excel over the NET CASH FLOW)